You want to send your child to one years (12 months) of boarding school starting 6 months from now. The cost of the school is $7,000 per month for 12 months. If your investments earn 3% APR, compounded monthly, how much would you have to invest today to cover the cost of the boarding school?
Answer #1
Annual Interest rate: 3%
Monthly Interest rate: (3/12) = 0.25%
Month Payments PVf at 0.25% Present value
6 7000 0.98513 6895.913
7 7000 0.982674 6878.716
8 7000 0.980223 6861.562
9 7000 0.977779 6844.451
10 7000 0.97534 6827.382
11 7000 0.972908 6810.356
12 7000 0.970482 6793.373
13 7000 0.968062 6776.432
14 7000 0.965648 6759.533
15 7000 0.963239 6742.676
16 7000 0.960837 6725.862
17 7000 0.958441 6709.089
Present value f payment made 81625.35
Amount required today: $ 81625.35
Note:
PVF os computed as under:
PVF for 6th mo;nth = 1 / (1.0025)^6 = 0.98513
PVF for 7th month = 1/ (1.0025)^7 = 0.982674
and so on.